OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. D(5, -2) m = 2/5

OpenStudy (anonymous):

-3?

OpenStudy (anonymous):

@paki wait paki, does standard form means (y-y1) = m(x-x1)

OpenStudy (paki):

yes

OpenStudy (anonymous):

alright thanks. ax+by=c

OpenStudy (anonymous):

oh, then no need to find be y+2 = 2/5 (x-5)

OpenStudy (anonymous):

b**

OpenStudy (anonymous):

thank you @paki

OpenStudy (anonymous):

2x + y = -2?

OpenStudy (anonymous):

just put it in standard form

OpenStudy (anonymous):

oh, if you simplify it. (2/5)x -y = 4

OpenStudy (anonymous):

and if you want to get rid of the fraction. that equation is also the same as 5 ( (2/5)x -y= 4 ) 2x -5y = 20

OpenStudy (anonymous):

The space it gives me is _x_ _y _ _ _ _ _ The answer has to fit in there

OpenStudy (anonymous):

each _ is a number value right?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

I don't think there can be a fraction though

OpenStudy (anonymous):

Thats why I'm so confused

OpenStudy (anonymous):

then how would you put in your old answer 2x+y= -2? You still have blank spaces. Have you tried typing in 2x -5y = 20 then hit enter?

OpenStudy (anonymous):

I think that's the only way, without fractions. You can use decimal places to fill up the space too.

OpenStudy (anonymous):

That might work actually I'll let you know if its right thank you